OpenStudy (anonymous):

If an associate’s average sale to 5 customers is $125, and her sales to four of the customers are $200, $150, $125, and $75, what is the associate’s sale to the fifth customer?

OpenStudy (anonymous):

summation of sales to those five customers=5*125=625 the sale to the fifth customer=625-200-150-125-75=75

OpenStudy (diyadiya):

let x be the sale to the fifth customer \[\frac{200+150+125+75+x}{5}=125\] \[\frac{550+x}{5}=125\] 550+x=625 x=625-550 x=75

OpenStudy (anonymous):

Thanks